# Surprisal and Metaphor Novelty: Moderate Correlations and Divergent Scaling Effects

### v2

@@ Abstract
Novel metaphor comprehension involves complex semantic processes and linguistic creativity, making it an interesting task for studying language models (LMs). This study investigates whether surprisal, a probabilistic measure of predictability in LMs, correlates with different metaphor novelty datasets. We analyse surprisal from 16 LM variants on corpus-based and synthetic metaphor novelty datasets. We explore a cloze-style surprisal method that conditions on full-sentence context. Results show that LMs yield significant moderate correlations with scores/labels of metaphor novelty. We further identify divergent scaling patterns: on corpus-based data, correlation strength decreases with model size (inverse scaling effect), whereas on synthetic data it increases (Quality–Power Hypothesis). We conclude that while surprisal can partially account for annotations of metaphor novelty, it remains a limited metric of linguistic creativity. [^1] Code and data are publicly available: https://github.com/OmarMomen14/surprisal-metaphor-novelty.
Surprisal and Metaphor Novelty: Moderate Correlations and Divergent Scaling Effects
@@ 1 Introduction
Recent advances in neural language modelling have led to a renewed interest in studying linguistic creativity, an aspect of language that was notoriously challenging for traditional NLP systems. A well-known instance of creative language is metaphor, which arises through mapping of a source domain–a lexical unit’s literal meaning–onto a target domain–its figurative meaning-Lakoff and Johnson (1980). Yet, not all metaphors are equally novel or creative as many of such mappings are highly conventionalized, such as in the famous example “to attack an argument,” where the mapping from the source domain “WAR” onto the target domain “ARGUMENT” corresponds to a conventional sense of the verb “attack” which can be found in dictionaries. Other metaphors, as in “The arrested water” (Figure 1) establish a more unconventional or novel mapping that is considered creative.
@@
However, separating novel metaphors from conventional ones remains a challenging task for human-based annotations Do Dinh et al. (2018); Reimann and Scheffler (2024). While computational methods on metaphor detection and understanding do not even take this distinction into account in most cases (Maudslay and Teufel, 2022), and some existing studies highlight that novel metaphors are more difficult to detect than conventional ones Neidlein et al. (2020); Tong et al. (2024). In this work, we propose to fill this gap and study questions on LMs processing of novel and conventional metaphors, and investigate LM-based metrics correlations to different setups of metaphor novelty annotations.
@@
However, recent work draws a mixed picture in terms of which LMs can provide the most robust and cognitively plausible predictors. Oh and Schuler (2023b) observe an inverse scaling effect when testing surprisal estimates from GPT-2 models of different sizes, showing that the surprisal computed with smaller model sizes achieved a better fit with human reading times than larger model sizes. Wilcox et al. (2023), on the other hand, train LMs of small and medium size on a range of languages and find that LM quality generally correlates with its psychometric predictive power.
In this paper, we investigate surprisal on metaphor words computed with LMs as a metric of metaphor novelty. We find significant moderate correlations between LMs’ surprisal and different metaphor novelty annotations. We perform our experiments on four different metaphor novelty datasets coming from corpus-based and synthetic setups, using LMs of different sizes and architectures. Most interestingly, we observe effects supporting the "inverse scale" pattern Oh and Schuler (2023b) on the two corpus-based novelty datasets, and contrary patterns supporting the counter-argument, "Quality-Power" hypothesis Wilcox et al. (2023) on the two synthetic datasets. We also investigate the effect of instruction-tuning on surprisal correlation to metaphor novelty. Moreover, we conduct a deeper analysis of the genre splits, revealing that genre, metaphor density and LM perplexity are potential factors underlying the quality of surprisal as a predictor of metaphor novelty. Finally, we introduce a new method of computing surprisal, cloze-surprisal, to include the right context of the word in its conditional probability. We find that this method can boost the correlation of surprisal with metaphor novelty annotations by a few points. In general, our study establishes a promising direction for studying linguistic creativity with LMs and calls for novel measures and datasets that provide systematic annotations of metaphor novelty across genres.
@@ 2.2 Predictive Powers of Surprisal
Surprisal estimates from LMs have been shown to capture human processing difficulty across multiple behavioural and neural measures. In self-paced reading and eye-tracking studies, token-level surprisals from causal models significantly predict reading times beyond lexical and syntactic factors Goodkind and Bicknell (2018); Wilcox et al. (2020); Oh and Schuler (2023a). Similar effects have been reported for acceptability judgments, where LM surprisals align with human sensitivity to gradient grammaticality Lau et al. (2017); Meister et al. (2021); Hu et al. (2020); Tjuatja et al. (2025).
@@ 2.3 LM-Based Metrics for Metaphor Novelty
Recent work has explored diverse LM-based approaches—prompting, surprisal, embedding similarity and attention patterns—for the analysis of metaphorical language. Aghazadeh et al. (2022) demonstrate that metaphor-relevant information is encoded in mid-layer embeddings of multilingual pre-trained LMs. Ichien et al. (2024) show that GPT-4 generates interpretations of novel literary metaphors that are favoured by human judges over interpretations by college students, suggesting sensitivity to metaphorical meaning beyond lexical overlap. Pedinotti et al. (2021) tested BERT’s masked token probabilities across conventional, novel, and nonsensical metaphors, showing that novel metaphors tend to receive lower probabilities, but distinctions between novelty and deviancy remained unclear. We extend this study towards more recent LMs and broader comparisons between models, datasets, and genres. Djokic et al. (2021) trained a BERT-based classifier to predict novelty scores jointly with the task of metaphor detection.
@@ 3 Metaphor Novelty Datasets
This Section describes the four datasets we use in our experiments and explains how metaphor novelty can be annotated in different approaches. Datasets statistics are reported in Tables 6 and 7.
@@ 3.2 Synthetic Datasets
Lai et al. (2009) investigated how our brains handle conventional and novel metaphors differently. To this end, they constructed a set of controlled experimental items featuring metaphors with two degrees of novelty. Two linguists selected 104 words, and constructed 4 sentences for each word, according to the Conceptual Metaphor Theory (CMT) Lakoff and Johnson (1980). For each word, the items include one (i) literal use, (i) conventional metaphor, (iii) novel metaphor and (iv) anomalous use, with the target word as the last word in each sentence. Familiarity and interpretability tests showed a significant difference between the conventional and novel metaphoric senses. For our experiments, we select only the conventional and novel metaphor senses for each word, resulting in 208 sentences (104 conventional and 104 novel metaphors).
@@ 4.1 Surprisal of a target word
In information theory Shannon (1948), the information content of an event $x$ with probability $p(x)$ is defined as: $I(x)=-\log p(x)$.
In the context of LMs trained to predict the next token in a sequence, this quantity is called surprisal and computed for a $w_{i}$ in a sequence such that: $\mathrm{Surprisal}(w_{i})=-\log p(w_{i}\mid w_{<i})$ [^3] We use $\log$ of base $e$ for all surprisal/perplexity computations in our study..
In our experiments, we measure word-level surprisal of metaphoric words in their sentence-level context. We feed every sentence to an LM in an independent teacher-forced forward pass, and record surprisal of the target (metaphoric) word(s). We denote this quantity as direct-surprisal to distinguish from cloze-surprisal.
Naturally, the method of computing word-level surprisal from LMs is a rough approximation process, and design choices of this method can bias the results Pimentel and Meister (2024); Lesci et al. (2025). We identify a couple of choices in the details of this method that can affect  [^4] We find differences in the correlation values when these choices are altered. the final results of our study. Mainly, subword tokenisation is the reason for such methodological confounds.
Firstly, the goal is to get the probability of a target word (e.g. “arrested” in Figure 1); however, most LMs produce probabilities per tokens (subwords) (e.g. [‘‘Ġarrest’’, ‘‘ed’’]) rather than words. And the method of aligning these tokens from the list of sentence tokens to specific words becomes a methodological choice. We choose to align the tokens to words by precomputing the offsets of the exact [^5] We make sure no punctuation or any other characters outside of the word boundary are included inside the offsets. target words in their corresponding sentences, and then searching for the minimal span of tokens in the LM’s tokenisation of the input sentence covering the target word’s offsets. This usually results in a proper alignment (e.g. ‘‘Ġarrested’’ for “arrested”), with very rare cases when the last token is attached to a punctuation (e.g. [‘‘Ġindivid’’, ‘‘ual’’, ‘‘ism,’’] for “individualism”). We also apply the corrections made by Pimentel and Meister (2024); Oh and Schuler (2024) that address the problem of leading whitespaces in tokenisations of most causal LMs.
Secondly, surprisal is computed using the conditional probability of the current word being the target word given the preceding words in the sentence. This makes computing surprisal of the first word in a sentence an issue [^6] In some cases in the VUA datasets, the target word is the first word in the sentence.. To allow computing surprisal of the first word in a sentence, we prepend the input sentence with a “begining of sequence” special token.
@@
We try to address this inherent weakness of word-level surprisal computed from causal models, which does not take the right context of the metaphoric word into account. This is an issue when sentences do not end in the metaphorical word, but feature various positions of the metaphorical word as in naturally occurring corpus data (see Figure 4), and also in some cases of the GPT-4o-metaphors dataset. To address this, we experiment with a new strategy for computing word-level surprisal autoregressively, while introducing the full sentence context: for each metaphor word, we input a prompt (as illustrated in Figure 1) that indicates a cloze test and show the whole sentence to the model except the metaphor word (replaced by a blank space), and then spell out the whole sentence again as if the model is solving the cloze test. We measure the surprisal of the target word at the second occurrence of the sentence, ensuring that the left and right context are included in the conditional probability of the metaphor word.
@@ 4.2 Evaluating the correlation
To determine whether LMs surprisal correlates with different metaphor novelty annotations, we measure multiple correlation metrics between surprisal and metaphor novelty scores/labels. For continuous novelty scores (as in VUA-ratings), we compute Pearson’s $r$ Pearson (1895) and Spearman’s $\rho$ Spearman (1904) correlations. For binary labels (conventional vs. novel), we compute the Rank-biserial $r_{b}$ Glass (1966) correlation. Rank-biserial estimates the probability that a random observation from the set of novel metaphors has a larger surprisal than one from the set of conventional metaphors, minus the reverse probability. We also estimate the potential of surprisal as a discriminator for binary novelty labels using the Area Under the ROC Curve (AUC) Fawcett (2006). We also compute the significance of all these estimates. All our metrics and tests do not assume normality, except for Pearson’s, for which we also provide a non-parametric alternative (Spearman).
@@ 4.3 Settings
Scales of surprisal can differ from one genre to another. Perplexity of a model on a certain dataset is the exponential of the average token-level surprisal of all tokens in the dataset. In that sense, perplexity indicates how much a certain model is “surprised” on average when predicting a certain dataset. We measure models’ perplexity on the genre splits of VUA by feeding sentences one by one to the model, measuring token-level surprisals accordingly and averaging and exponentiating them to obtain perplexity. This yields higher values of perplexity than common values reported in literature, as we use a shorter context (single sentences).
@@ 4.4 Summary
Our experiments rely on a large collection of surprisal measures for each metaphor word in the four datasets: 32 different surprisal values (including direct and cloze) for 15,155 metaphoric words of the VUA dataset and for 208 and 200 metaphoric words of Lai2009 and GPT-4o-metaphors, respectively. In Figure 3, we plot the distributions of surprisal values by model, in addition to metaphor novelty scores/labels from VUA. From these plots and further normal distribution tests, we find that novelty scores and surprisal values are not strictly normally distributed.
@@ 5.1 Overall Correlation Results
Generally, we find LMs’ direct surprisal values correlate positively with metaphor novelty annotations across the four datasets (Tables 1 and 2). All correlation estimates are statistically significant for both the corpus-based and synthetic datasets  [^7] The relatively large corpus-based datasets (15,155 datapoints) can bias the significance values; however, we get similar significance values for the relatively small synthetic datasets (100 datapoints).. Overall for direct surprisal, the largest Pearson correlation $r=.419$, largest Spearman correlation $\rho=.417$, largest rank-biserial correlation $r_{b}=.638$ and largest $AUC=.819$ come from GPT2-base on VUA-ratings.
We also find that the correlation strengths’ scales differ across the different annotation datasets. By comparing the rank-biserial estimate $r_{b}$ across the four datasets, we find its ranges as follows: (.47-.64) for VUA-ratings, (.43-.58) for VUA-dictionary, (.28-.50) for Lai2009 and (.38-.63) for GPT-4o-metaphors. These results show that surprisal correlates relatively more with the corpus-based datasets than the synthetic datasets. Also, surprisal correlates relatively more with human ratings on VUA than the dictionary-based annotation approach. While the most controlled dataset (Lai2009) gets the weakest correlations with surprisal.
@@ 5.2 Model Size Effects
The effect of model size is consistent (but diverging in direction) across the two dataset types. We plot the rank-biserial correlations of the ten model variants for the four datasets in Figure 2. The two corpus-based datasets (blue) show a monotonically decreasing correlation as model size increases per model family. On the other hand, the two synthetic datasets (red) exhibit an increasing correlation behaviour as model size increases, with a single minor exception at the GPT2 model family on the GPT-4o-metaphors dataset.

### v3

# Surprisal and Metaphor Novelty Judgments: Moderate Correlations and Divergent Scaling Effects Revealed by Corpus-Based and Synthetic Datasets

@@ Abstract
Novel metaphor comprehension involves complex semantic processes and linguistic creativity, making it an interesting task for studying language models (LMs). This study investigates whether surprisal, a probabilistic measure of predictability in LMs, correlates with annotations of metaphor novelty in different datasets. We analyse the surprisal of metaphoric words in corpus-based and synthetic metaphor datasets using 16 causal LM variants. We propose a cloze-style surprisal method that conditions on full-sentence context. Results show that LM surprisal yields significant moderate correlations with scores/labels of metaphor novelty. We further identify divergent scaling patterns: on corpus-based data, correlation strength decreases with model size (inverse scaling effect), whereas on synthetic data it increases (quality–power hypothesis). We conclude that while surprisal can partially account for annotations of metaphor novelty, it remains limited as a metric of linguistic creativity. [^1] Code and data are publicly available: https://github.com/OmarMomen14/surprisal-metaphor-novelty.
Surprisal and Metaphor Novelty Judgments: Moderate Correlations and Divergent Scaling Effects Revealed by Corpus-Based and Synthetic Datasets
@@ 1 Introduction
Recent advances in language modelling have led to a renewed interest in studying linguistic creativity, an aspect of language that was notoriously challenging for traditional NLP systems. A well-known instance of creative language is metaphor, which arises through mapping of a source domain–a lexical unit’s literal meaning–onto a target domain–its figurative meaning-Lakoff and Johnson (1980). Yet, not all metaphors are equally novel or creative as many of such mappings are highly conventionalized, such as in the famous example “to attack an argument,” where the mapping from the source domain “WAR” onto the target domain “ARGUMENT” corresponds to a conventional sense of the verb “attack” which can be found in dictionaries. Other metaphors, as in “The arrested water” (Figure 1) establish a more unconventional or novel mapping that is considered creative.
@@
However, separating novel metaphors from conventional ones remains a challenging task for human annotators Do Dinh et al. (2018); Reimann and Scheffler (2024). While computational methods on metaphor analysis do not often take the novelty dimension into account (Maudslay and Teufel, 2022), and some existing studies highlight that novel metaphors are more difficult to detect than conventional ones Neidlein et al. (2020); Tong et al. (2024). In this work, we propose to fill this gap by studying questions related to LMs’ processing of novel and conventional metaphors and investigating correlations between LM-based metrics and different setups of metaphor novelty annotations.
@@
However, recent work draws a mixed picture in terms of which LMs can provide the most robust and cognitively plausible predictors for processing difficulty. Oh and Schuler (2023b) observe an inverse scaling effect when testing surprisal estimates from GPT-2 models of different sizes, showing that surprisal computed with smaller model sizes achieved a better fit with human reading times than larger models. Wilcox et al. (2023), on the other hand, train LMs of small and medium size on a range of languages and find that LM quality generally correlates with its predictive power of reading times.
In this paper, we investigate surprisal of metaphoric words computed using a selection of LMs as a metric of metaphor novelty. We find significant, positive and moderate correlations between LMs’ surprisal and different metaphor novelty annotations. We perform our experiments on four different datasets coming from corpus-based and synthetic setups, using LMs of different sizes. Most interestingly, we observe effects supporting the "inverse scale" pattern Oh and Schuler (2023b) on the two corpus-based novelty datasets, and contrary patterns supporting the counter-argument, "Quality-Power" hypothesis Wilcox et al. (2023) on the two synthetic datasets. We also investigate the effect of instruction-tuning on surprisal correlation to metaphor novelty. Moreover, we conduct a deeper analysis of the genre splits, revealing that genre, metaphor density, and LM perplexity are potential factors underlying the quality of surprisal as a predictor of metaphor novelty. Finally, we introduce a new method of computing surprisal, cloze-surprisal, to include the right context of the word in its conditional probability. We find that this method can boost the correlation of surprisal with metaphor novelty annotations by a few points. In general, our study establishes a promising direction for studying linguistic creativity with LMs and calls for novel measures and datasets that provide systematic annotations of metaphor novelty across genres.
@@ 2.2 Predictive Powers of Surprisal
Surprisal estimates from LMs have been shown to capture human processing difficulty across multiple behavioural and neural measures. In self-paced reading and eye-tracking studies, token-level surprisals from causal models significantly predict reading times beyond lexical and syntactic factors Goodkind and Bicknell (2018); Wilcox et al. (2020); Oh and Schuler (2023a). Similar effects have been reported for acceptability judgments, where LM surprisals align with human sensitivity to grammaticality Lau et al. (2017); Meister et al. (2021); Hu et al. (2020); Tjuatja et al. (2025).
@@ 2.3 LM-Based Metrics for Metaphor Novelty
Recent work has explored diverse LM-based approaches—prompting, surprisal, embedding similarity, and attention patterns—for the analysis of metaphorical language. Aghazadeh et al. (2022) demonstrate that metaphor-relevant information is encoded in mid-layer embeddings of multilingual pre-trained LMs. Ichien et al. (2024) show that GPT-4 generates interpretations of novel literary metaphors that are favoured by human judges over interpretations by college students, suggesting sensitivity to metaphorical meaning beyond lexical overlap. Pedinotti et al. (2021) tested BERT’s masked token probabilities across conventional, novel, and nonsensical metaphors, showing that novel metaphors tend to receive lower probabilities, but distinctions between novelty and nonsense remained unclear. We extend this study towards more recent LMs and broader comparisons between models, datasets, and genres. Djokic et al. (2021) trained a BERT-based classifier to predict novelty scores jointly with the task of metaphor detection.
@@ 3 Metaphor Novelty Datasets
This Section describes the four datasets we use in our experiments and explains how metaphor novelty is annotated in different approaches. Datasets statistics are reported in Tables 6 and 7 (Appendix A.1).
@@ 3.2 Synthetic Datasets
Lai et al. (2009) investigated how our brains handle conventional and novel metaphors differently. To this end, they constructed a set of controlled experimental items featuring metaphors with two degrees of novelty. Two linguists selected 104 words and constructed 4 sentences for each word, according to the Conceptual Metaphor Theory (CMT) Lakoff and Johnson (1980). For each word, the items include one (i) literal use, (i) conventional metaphor, (iii) novel metaphor, and (iv) anomalous use, with the target word as the last word in each sentence. Familiarity and interpretability tests showed a significant difference between the conventional and novel metaphoric senses. For our experiments, we select only the conventional and novel metaphor senses for each word, resulting in 208 sentences (104 conventional and 104 novel metaphors).
@@ 4.1 Surprisal of a target word
In information theory Shannon (1948), the information content of an event $x$ with probability $p(x)$ is defined as: $I(x)=-\log p(x)$. In the context of LMs trained to predict the next token in a sequence, this quantity is called surprisal and computed for a $w_{i}$ in a sequence such that: $\mathrm{Surprisal}(w_{i})=-\log p(w_{i}\mid w_{<i})$ [^3] We use $\log$ of base $e$ for all surprisal/perplexity computations in our study..
In our experiments, we measure word-level surprisal of metaphoric words in their sentence-level context. We feed every sentence to an LM in an independent teacher-forced forward pass, and record surprisal of the target (metaphoric) word(s). We denote this quantity as direct-surprisal to distinguish it from cloze-surprisal.
Most LMs operate on subwords rather than words, so deriving word-level probabilities requires implementation choices that can affect surprisal estimates and, in turn, downstream measures Pimentel and Meister (2024); Lesci et al. (2025). For transparency and reproducibility, we report two choices in our implementation that affect the computed surprisal values. First, we choose to align tokens (subwords) to words by precomputing the offsets of the exact [^4] We make sure no punctuation or any other characters outside of the word boundary are included inside the offsets. target words in their corresponding sentences, and then searching for the minimal span of tokens in the LM’s tokenisation of the input sentence covering the target word’s offsets. This usually results in a proper alignment with the target word in addition to a leading whitespace character (e.g. ‘‘Ġarrested’’ for “arrested”), with very rare cases when the last token is attached to a punctuation (e.g. [‘‘Ġindivid’’, ‘‘ual’’, ‘‘ism,’’] for “individualism”). We also apply the corrections made by Pimentel and Meister (2024); Oh and Schuler (2024) that address the problem of leading whitespaces in tokenisations of most causal LMs. Second, surprisal is computed using the conditional probability of the current word being the target word given the preceding words in a sentence. This makes computing surprisal of the first word in a sentence an issue [^5] In some cases in the VUA datasets, the target word is the first word in the sentence.. To compute surprisal of the first word in a sentence, we prepend the input sentence with a "beginning of sequence" special token.
@@
A key limitation of standard (left-to-right) surprisal as a proxy for semantic contextual properties (e.g., metaphor novelty) is that it does not condition on future (right) context in the sentence. This matters in naturally occurring corpus data (e.g VUA datasets), where metaphorical words can appear in many sentence positions, and in some cases in the GPT-4o-metaphors dataset (see Table 8). To incorporate the right context while retaining an autoregressive setup, we compute cloze-surprisal: for each target word, we prompt the model to predict a missing word in a sentence, and we replace the target in the sentence by a blank space, and then we append the same sentence again as the intended completion (Figure 1). We measure surprisal for the target word at its position in the second occurrence of the sentence. Because the first occurrence exposes the full right context, the resulting conditional probability effectively incorporates both left and right contexts of the target word.
@@ 4.2 Evaluating the correlation
To determine whether LMs’ surprisal correlates with different metaphor novelty annotations, we use multiple correlation metrics between surprisal and metaphor novelty scores/labels. For continuous novelty scores (as in VUA-ratings), we compute Pearson’s $r$ Pearson (1895) and Spearman’s $\rho$ Spearman (1904) correlations. For binary labels (conventional vs. novel), we compute the Rank-biserial $r_{b}$ Glass (1966) correlation. Rank-biserial estimates the probability that a random observation from the set of novel metaphors has a larger surprisal than one from the set of conventional metaphors, minus the reverse probability. We also estimate the potential of surprisal as a discriminator for binary novelty labels using the Area Under the ROC Curve (AUC) Fawcett (2006). We also compute the significance of all these estimates. All our metrics and tests do not assume normality, except for Pearson’s, for which we also provide a non-parametric alternative (Spearman).
@@ 4.3 Settings
Scales of surprisal can differ from one genre to another, and from one model to another. To this end, we investigate perplexity as a potential factor in our study. Perplexity of a model on a certain dataset is the exponential of the average token-level surprisal of all tokens in the dataset. In that sense, perplexity indicates how much a certain model is “surprised” on average when predicting a certain dataset. We measure models’ perplexity on the genre splits of VUA by feeding sentences one by one to the model, measuring token-level surprisals accordingly and averaging and exponentiating them to obtain perplexity. This yields higher values of perplexity than common values reported in literature, as we use a shorter context (single sentences).
@@ 4.4 Summary
Our experiments rely on a large collection of surprisal measures for each metaphor word in the four datasets: 32 different surprisal values (including direct and cloze) for 15,155 metaphoric words of the VUA dataset and for 208 and 200 metaphoric words of Lai2009 and GPT-4o-metaphors, respectively. In Figure 3 (Appendix A.6), we plot the distributions of surprisal values by model, in addition to metaphor novelty scores/labels from VUA. From these plots and further normal distribution tests, we find that novelty scores and surprisal values are not strictly normally distributed.
@@ 5.1 Overall Correlation Results
Generally, we find LMs’ direct surprisal values correlate positively with metaphor novelty annotations across the four datasets (Tables 1 and 2). All correlation estimates are statistically significant for both the corpus-based and synthetic datasets  [^6] The relatively large corpus-based datasets (15,155 datapoints) can bias the significance values; however, we get similar significance values for the relatively small synthetic datasets (100 datapoints).. Overall, for direct surprisal, the largest Pearson correlation $r=.419$, largest Spearman correlation $\rho=.417$, largest rank-biserial correlation $r_{b}=.638$ and largest $AUC=.819$ come from GPT2-base on VUA-ratings. These values indicate a significant positive correlation, yet its strength is moderate.
We also find that the correlation strengths’ ranges differ across the different annotation datasets. By comparing the rank-biserial estimate $r_{b}$ across the four datasets, we find its ranges as follows: (.47-.64) for VUA-ratings, (.43-.58) for VUA-dictionary, (.28-.50) for Lai2009, and (.38-.63) for GPT-4o-metaphors. These results show that surprisal correlates more strongly with the corpus-based data than with the synthetic data. Also, surprisal correlates more strongly with human ratings on VUA than the dictionary-based annotation approach. While the most controlled dataset (Lai2009) gets the weakest correlations with surprisal.
@@ 5.2 Model Size Effects
The effect of model size is consistent (but diverging in direction) across the two dataset types. We plot the rank-biserial correlations of the ten model variants for the four datasets in Figure 2. For the two corpus-based datasets (blue), surprisal–novelty correlation is monotonically decreasing as model size increases (per model family). On the other hand, for the two synthetic datasets (red), surprisal–novelty correlation increases as model size increases, with a single minor exception at the GPT2 model family on the GPT-4o-metaphors dataset.